The average EV/EBITDA ratio of the competitor firms to Jorge's Junk Removal is 6.5. Jorge's Junk Removal has an EBIT of $400 million, depreciation of $100 million, no amortization expense, $500 million in debt, $40 million in cash, and 25 million outstanding shares.   What should the price of a share of Jorge's Junk Removal be?

A) $59.60

B) $131.60

C) $32.50

D) $85.60

E) $111.60
Answer #1

EBIDTA = EBIT+depr = 400+100 = 500

EV = EBIDTA*EV/EBIDTA = 500*6.5 = 3250

Enterprise value = Equity value+ MV of debt
- Cash & Cash Equivalents
3250 = Equity value+500-40
Equity value = 2790
share price = equity value/number of shares
share price = 2790/25
share price = 111.6